If 0.1 g of Ferric sulfate Fe_2(SO_4)_3 is added t
Answer #1

1) Ferric Sulphate Mol wt = 400

Mass of ferric sulfate = 0.1 gms

No of moles = 0.1/400 = 0.00025

Therefore Molarity = 0.00025 M

100 mg/L or 100 ppm

0.01%

2) PV=nRT

V= 10 L, T= 293.15K, n= 500/70.906 (7.05 moles), R=0.0821 atm L mol-1 K-1

P = 7.05*0.0821*293.15/10

P= 16.97 atm

3) V1/T1 = V2/T2 (n,p constant)

10000/283.15 = X/300.15

Therefore X = 10600.39 m3

5) a) E.N of B = 2

E.N of C = 2.5

E.N of Cl = 3

Since E.N difference between B-Cl is more, it is more polar

b) E.N of P = 2.2

E.N of F = 4

E.N of Cl = 3

Since E.N difference in P-F is more it is more polar